OpenStudy (anonymous):

the system of equations is coincident. What are the missing values? 3x+y=7 9x+__Y=__

OpenStudy (anonymous):

@satellite73 @amistre64 @lsherron

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@RyleeC

OpenStudy (amistre64):

what defines a set of coincident lines?

OpenStudy (anonymous):

what do you mean

OpenStudy (amistre64):

its your question, not mine .... "the system of equations is coincident" define coincident lines for me

OpenStudy (anonymous):

Two lines or shapes that lie exactly on top of each other.

OpenStudy (amistre64):

good, now in this case, when one equation is a scaled version of the other, they are coincident: x + y = c and n(x + y = c)

OpenStudy (amistre64):

3x + y = 7 is coincident to: 3nx + ny + 7n but we are given that 3nx = 9x, so what is n?

OpenStudy (amistre64):

* = 7n typoed that little bit